the titration of 19.30 mL of HBr solution of unknown concentration requires 18.11 mL of a 0.100 M KOH solution to reach the equivalence point. What is the concentration of the unknown HBr solution in M?
1
Expert's answer
2020-04-13T02:32:12-0400

In equivaelence point n(HBrinitial)=n(KOHadded). Then if concentration of HBR was x,

19.30/1000*x = 18.11/1000*0.1

"1={\frac {18.11*0.100} {19.30*x}}"

"x={\frac {18.11*0.100} {19.30}}=0.094"

Concentration of HBr solution is 0.094 M